OpenStudy (anonymous):

An auditorium earned $30,000 in sold-out concert ticket sales. Front section tickets cost $75 per seat and back section tickets cost $50 per seat. The number of front section seats is twice the number of back section seats. How many seats are in the front section?

OpenStudy (anonymous):

30,000=75x+50y x is front seats and y is back seats x=2y

OpenStudy (anonymous):

150*50=7500 300*75=22500 30,000=75(300)+50(150)